# I2VControl-Camera: Precise Video Camera Control with Adjustable Motion Strength

### v1

@@ Abstract
Video generation technologies are developing rapidly and have broad potential applications. Among these technologies, camera control is crucial for generating professional-quality videos that accurately meet user expectations. However, existing camera control methods still suffer from several limitations, including control precision and the neglect of the control for subject motion dynamics. In this work, we propose I2VControl-Camera, a novel camera control method that significantly enhances controllability while providing adjustability over the strength of subject motion. To improve control precision, we employ point trajectory in the camera coordinate system instead of only extrinsic matrix information as our control signal. To accurately control and adjust the strength of subject motion, we explicitly model the higher-order components of the video trajectory expansion, not merely the linear terms, and design an operator that effectively represents the motion strength. We use an adapter architecture that is independent of the base model structure. Experiments on static and dynamic scenes show that our framework outperformances previous methods both quantitatively and qualitatively. $\dagger$ Corresponding author
@@ 1 Introduction
Video generation technologies are explored to synthesize dynamic and coherent visual content, conditioned on various modalities including text (Blattmann et al., 2023c; Wang et al., 2024a; Gupta et al., 2023) and images (Blattmann et al., 2023b; Chen et al., 2024). Video generation has broad application potential across various fields, such as entertainment, social media, and film production. Motion controllability is crucial for ensuring that generated videos accurately meet user expectations, with camera control being one of the most important aspects. Camera control is the process of adjusting the position, angle, and motion of a camera, resulting in changes to the composition, perspective, and dynamic effects of a video. This technique is essential for generating professional-quality videos, as it influences the attention of viewers and enhances the expressiveness of scenes.
@@
Figure 1: We propose I2VControl-Camera, a novel camera control method for image-to-video generation. Our approach supports any camera movement style, offering high control precision and stability, ensuring natural dynamics and adjustable motion strength, which can be seen in Sec. 4.
@@ 3.2 Control Signal Construction
Figure 2: Illustration of point trajectory function. The point cloud lifted from 2D to 3D are originally in the world coordinate system, and then transformed to the initial camera coordinate system. In this way, the camera movement is converted as point trajectorys.
@@ 3.3 Data Pipeline
To address the first issue, we employ metric depth estimation method, Unidepth (Piccinelli et al., 2024), to bridge the gap between 2D and 3D data representation. For the second issue, we utilize a tracking method, Cotracker (Karaev et al., 2023), to establish pixel correspondence between consecutive frames, so that we can obtain the discrete trajectory (still denoted as $\mathcal{F}(\mathbf{p},\lambda)$ for ease of reading). For the third issue, we need to extract $\Omega_{S},\Omega_{D}\subseteq\Omega$ from $\Omega$. A key insight lies in Eq. 10, which means the trajectory on $\Omega_{S}$ can be linearly fitted well. We solve this problem in a iterative manner, as described in Alg. 1.
@@ 3.4 Network, Training and Inference
During the training phase, we merely incorporate the insertion of the control signal, while all other training strategies remain unchanged. We adopt the same loss function and the same scheduler, with the sole modification being the introduction of the control signal condition. During testing, we first obtain the metric depth of the input image using Unidepth (Piccinelli et al., 2024), then compute the trajectory based on the input camera motion parameters (as in Fig. 2), and require users to input the motion strength. As a result, this conforms to the training paradigm and produces suitable camera movement effects. When the motion strength is set small, nearly static camera movements can be achieved, whereas a significant motion strength allows for more pronounced dynamics of the subject.
@@ 4.2 Visualization Results
In this section, we show the visualization results of our method, demonstrating both pixel-level controllability and the motion strength adjustment. Due to the length and format of the paper, the results shown below are all in frame-by-frame image format.
@@ 5 Conclusion
Acknowledgement
We thank Yuxi Xiao for his generous help and advice with the tracking algorithm.

### v3

@@ Abstract
Video generation technologies are developing rapidly and have broad potential applications. Among these technologies, camera control is crucial for generating professional-quality videos that accurately meet user expectations. However, existing camera control methods still suffer from several limitations, including control precision and the neglect of the control for subject motion dynamics. In this work, we propose I2VControl-Camera, a novel camera control method that significantly enhances controllability while providing adjustability over the strength of subject motion. To improve control precision, we employ point trajectory in the camera coordinate system instead of only extrinsic matrix information as our control signal. To accurately control and adjust the strength of subject motion, we explicitly model the higher-order components of the video trajectory expansion, not merely the linear terms, and design an operator that effectively represents the motion strength. We use an adapter architecture that is independent of the base model structure. Experiments on static and dynamic scenes show that our framework outperformances previous methods both quantitatively and qualitatively. Project page: https://wanquanf.github.io/I2VControlCamera.
@@ 1 Introduction
Video generation technologies are explored to synthesize dynamic and coherent visual content, conditioned on various modalities including text (Blattmann et al., 2023c; Wang et al., 2024a; Gupta et al., 2023) and images (Blattmann et al., 2023b; Feng et al., 2024). Video generation has broad application potential across various fields, such as entertainment, social media, and film production. Motion controllability is crucial for ensuring that generated videos accurately meet user expectations, with camera control being one of the most important aspects. Camera control is the process of adjusting the position, angle, and motion of a camera, resulting in changes to the composition, perspective, and dynamic effects of a video. This technique is essential for generating professional-quality videos, as it influences the attention of viewers and enhances the expressiveness of scenes.
@@
Figure 1: We propose I2VControl-Camera, a novel camera control method for image-to-video generation, offering high control precision and adjustable motion strength.
@@ 3.2 Control Signal Construction
Figure 2: We lift the input image from 2D to 3D as a RGBD point cloud. When the camera moves, the 3D points can be considered as moving in the camera coordinate system. Then we project them onto 2D according to current camera pose to obtain the 2D point trajectory.
@@ 3.3 Data Pipeline
To address the first issue, we employ metric depth estimation method, Unidepth (Piccinelli et al., 2024), to bridge the gap between 2D and 3D data representation. For the second issue, we utilize a tracking method, SpatialTracker (Xiao et al., 2024), to establish pixel correspondence between consecutive frames, so that we can obtain the discrete trajectory (still denoted as $\mathcal{F}(\mathbf{p},\lambda)$ for ease of reading). For the third issue, we need to extract $\Omega_{S},\Omega_{D}\subseteq\Omega$ from $\Omega$. A key insight lies in Eq. 10, which means the trajectory on $\Omega_{S}$ can be linearly fitted well. We solve this problem in a iterative manner, as described in Alg. 1.
@@ 3.4 Network, Training and Inference
During the training phase, we merely incorporate the insertion of the control signal, while all other training strategies remain unchanged. We adopt the same loss function and the same scheduler, with the sole modification being the introduction of the control signal condition. During testing, we first employ Unidepth (Piccinelli et al., 2024) to lift the input image into a RGBD point cloud. When the user moves the camera, we convert it as the transformation of 3D points in the camera coordinate systemcan according to the camera poses. We finally project the transformed 3D points of each frame onto camera plane to compute the 2D trajectory, as shown in Fig. 2. Additionally, the user can provide a scalar value for the motion strength control. As a result, this conforms to the training paradigm and produces suitable camera movement effects. When the motion strength is set small, nearly static camera movements can be achieved, whereas a significant motion strength allows for more pronounced dynamics of the subject.
@@ 4.2 Visualization Results
In this section, we show the visualization results of our method, demonstrating both pixel-level controllability and the motion strength adjustment. Due to the format of the paper, the results shown below are all in frame-by-frame image format. Please see our project page for video results.